The power needed to operate your body is about 100 watts. Suppose a human body (of typical mass) could run on fusion power and could convert 0.8 % of its mass into energy.

For how long could this body then operate on fusion power?
Answer #1

mass = m = 60 kg

0.8% of mass = 0.8 x 60 / 100 = 0.48 kg

energy = E = m c^2 = 0.48 * ( 3 x 10^8 ) ^2 = 4.32 x 10^ 16 joules

power = P = E / t

time = t = E / P = 4.32 X 10^16 / 100 = 4.32 X 10^14 sec = 4.32 x 10^14 / 31556952 = 1.37 x 10^7 years